Commissioning Notes for 20080426 UT

Summary:

This will be the Friday night log which is tonight and thus has not happened yet. If you want the Thursday night log go to 24-25 Apr 2008

Running Notes/Log (UTtimes):

01:23 Open, ~40 minutes to sunset. Skyflats:
  • 012329 + 00001_cube (100x 0.01s, H filter) ~3000DN
  • 012504 + 00002_cube (-9.0 0.0 from 1st)
  • 012604 + 00003_cube (9.0 0.0 from 1st)
  • 012642 + 00004_cube (0.0 -9.0 from 1st)
  • 012713 + 00005_cube (0.0 9.0 from 1st) mean=1893. dark-subtracted counts.

  • 012921 + 00006_cube (100x 0.01s, J filter) ~2300DN
  • 013012 + 00007_cube ((-9.0 0.0 from 1st)
  • 013111 + 00008_cube (9.0 0.0 from 1st)
  • 013151 + 00009_cube (0.0 -9.0 from 1st)
  • 013240 + 00010_cube (0.0 9.0 from 1st)

Repeat J flats, longer exposures
  • 020203 + 00012_cube (no dithering, 25x0.25s)
  • 020236 + 00013_cube (no dithering, 25x0.25s)
  • 020347 + 00014_cube (no dithering, 25x0.25s)
  • 020424 + 00015_cube (no dithering, 25x0.25s)
  • 020440 + 00016_cube (no dithering, 25x0.25s) mean=3381. dark-subtracted counts

Calibration files are now available in IRTC$/Calib under LBTtools:
  • dark245K_2s.fits
  • dark245K_1s.fits
  • dark245K_500ms.fits
  • dark245K_250ms.fits
  • dark245K_100ms.fits
  • dark245K_50ms.fits
  • dark245K_10ms.fits
  • dark245K_1ms.fits

  • nHflat.fits
  • nJflat.fits

02:17 TCS install in progress...

03:00 Changed the rotator fixed angular offset from 199.1, which puts N-up on the IRTC GUI, to 19.1 to orient N-up on the images as displayed by IRAF. Checking...confirmed. On images displayed from iraf to DS9, at PA=0, N is now up and E is to the left.

03:39 pointto WT10_307. Comes up at the edge of the detector to the SE. Saturates in 10ms guider images.

03:44 pointto BS9148. GCS not acquiring the initial image, but it can be done by hand. IRTC GetParameter is not populating the image headers. Shutting down to push a new TCS build up. Problems took some time to resolve.

06:00 Went back to orientation with N down and E to the right as displayed on ds9. GCS was not able to guide when N was up and E left.

06:10 Guiding on-axis again. We are trying to find the right combination of pointing kernel parameters to allow correct orientation of the IRTC and correct guiding. Nope. Again, going back to what worked yesterday.

06:30 Walking the star manually over to the WFS aperture by updating the hotspot on the fly. This works, but it is slow. Exposure times on WFS and AG appear to be hard-wired to be the same?

06:48 Ran into the -1 degree limit on the rotator. Moving to a new source (WT10_354)

090849 wfsc000013 15s lots of astig. on BS9169 (Persson IR standard)

Clearing the active optics caused the guide star to walk off the guide box too fast for GCS to recover. Need to implement pointing-free coma corrections ASAP.

Some on-axis WFS data:
  • 092622 wfsc000015 Stage: 7.455 600.403
  • 092913 wfsc000016 Z6 -1000
  • 093156 wfsc000017 Z6 -500 This is a guider sub-image!
  • 093327 wfsc000018 Z6 -500
  • 093502 wfsc000019 Z6 0
  • 093630 wfsc000020 Z6 +500
  • 093840 wfsc000020 Z6 +1000 returned to zero when done.
there is another guider image in there, the last good wfsc here was 22.

Coma: (steps of 250 to keep guiding working)
  • 094216 wfsc000023 Z7 +250
  • 094346 wfsc000023 Z7 +500 the image capture command looks like it hung
  • 094559 wfsc000024 Z7 +750
  • 094704 wfsc000025 Z7 +1000

Calibrating RX RY vs. guider image coordinates:
  • Stages 7.412 600.399
  • /Repository/nom_ctr.fits
  • /Repository/rxminus4.fits RX -4.0 from nom_ctr (101537 for orientation)
  • /Repository/rxplus2.fits RX +2.0 from nom_ctr (101722)
  • /Repository/ryplus2.fits RY +2.0 from nom_ctr (101921)
  • /Repository/ryminus2.fits RY -2.0 from nom_ctr (102036)
  • /Repository/back2ctr.fits (102214) back on the nominal center starting position

10:28 More fan tests... T=4.3 WS2.24@az200 TelAZ86, TelEL=77
  • 104230 00029_cube on a star prior to tests, all systems running (1000 x 5ms)
  • 105337 00030_cube lower treehouse fans turned down, uppers still going strong
  • 110739 00031_cube only upper left treehouse fan turned down (others still going strong)
  • 112252 00032_cube both upper treehouse fans turned down (still vibrating, hmmm)
  • 113612 00033_cube all 4 fans slowed down.
  • 113824 00034_cube all 4 fans slowed down (they will speed up at some point).
  • 114040 00035_cube
  • 114325 00036_cube
  • 114553 00037_cube seems to be vibrating more now, about the time the fans should have sped up (~12 minutes).
  • 114926 00038_cube definitely back up to full shaking strength.

-- DavidThompson & NormCushing - 26 Apr 2008
Topic revision: r11 - 26 Apr 2008, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback